How is it written in a probability form, this question. If a family wants 3 kids, what are their chances of having at least one boy?
Answer by stanbon(75887)

How is it written in a probability form, this question. If a family wants 3 kids, what are their chances of having at least one boy?
---
P(at least one boy)
= 1 - P(3 girls)
= 1 - (1/2)^3
= 1-(1/8)
= 7/8
----------------
Cheers,
Stan H.